# PROSPECT: Precision Robot Spectroscopy Exploration and Characterization Tool

### v1

@@ V-C Surface Normal Matching
To orient the fiber optic normal to the surface at a fixed depth, a surface normalization algorithm was implemented. Given the initial pose of the Stewart platform, the positioning and readings of three time of flight sensors onboard the end effector, and a target offset distance, a new pose for the platform is computed. The platform is then commanded to move to this position using the joint angles computed by the inverse kinematics algorithm.
@@ VI Spectral-Spatial Modeling
Despite their slender form, fiber optic cables acquire light from the environment in a more complex manner than tracing a ray from the fiber terminus to the surface. We begin by modeling the interaction of light reflected from a target object and into the ferrule of the fiber optic cable. Numerical aperture (NA) defines the maximum angle from which light will be accepted into the fiber and is given by: $NA=n_{0}\sin{\theta_{max}}$. $n_{0}$ defines the refractive index of the medium through which light transits before entering the fiber, which we assume to be air with a refractive index of 1.0. $\theta_{max}=\arcsin{NA}$ yields the half-angle forming the acceptance cone.
Fig. 4 shows the fiber optic cable above a nominal surface $S$ that the PROSPECT platform will scan. The acceptance cone is centered on the ferrule. The distance from the tip of the ferrule to the surface is given as $d$, which intersects with the surface in a circle with a radius $r$. $d$ is either approximated from a static 3D scan of the object surface and confirmed using the 3 ToF distance sensors. The size of the cone base varies as a function of $d$, but $\theta_{max}$ remains constant. The area of the circular intersection region is given as $A=\pi(d\tan{(\arcsin{(NA)}}))^{2}$.
Figure 4: Model of spectroscopic measurement and modeling of object reflectance. Light is emitted from the PROSPECT platform (yellow) and reflected by the surface $S$ (green). The conical acceptance profile is defined by offset distance from the surface $d$, subsurface penetration $\epsilon$, and acceptance angle $\theta_{max}$. The intersection of the acceptance cone with the point cloud yields a subset of points $S_{c}$ with which the observed spectral signature is associated.
@@
To mark points as having been explored, we find the intersection between the surface $S$ and the circular profile of the acceptance cone projected on the surface, $S_{c}$. The intersection is found by creating a 3D mesh of the acceptance cone projected to slightly below the surface of $S$, at a distance $\epsilon$. The mesh is then transformed into Delaunay triangulation, which may be queried to find points from $S$ that lie inside it, $S_{c}$. Since scanning occurs over the surface at a close distance, the number of points observed follows the following relationship: $1\leq||S_{c}||<<||S||$. As the end-effector is translated according to Algorithm 1, the point cloud is annotated such that each point $x_{i},y_{i},z_{i}$ in $S_{c}$ is associated with a spectral vector $[\lambda_{0},\lambda_{1},...,\lambda_{n}]$.
As the size of $A$ increases, the intersecting area and points in $S_{c}$ also increases. However, as the distance $d$ increases, the number of points contributing to the spectral signal increases while the overall intensity of the signal decreases because fewer photons enter the fiber probe. This mixing of spectral signatures from multiple points complicates the interpretation and classification of the signal. Knowing the approximate point density of point cloud $S$, we chose a constant value for $d$ to avoid variable signal quality. Algorithm 2 details the steps to extract the points $S_{c}$ and form the correlated spectral-spatial measurements.
@@ VII-A Sparse Point Acquisition
The first set of experiments demonstrates how the creation of a sparse spectral-spatial point cloud. These scans are designed to be completed in less than 5 minutes to broadly understand an object’s surface composition. Fig.  5 demonstrates how spectral signature can vary throughout an object’s surface, emphasizing the importance of collecting precise spectral measurements across the entirety of the surface, shown for increasingly complex geometries. In each of these scenarios, viewpoints were planned with a voxel-size (point density) of 5 cm.
We test our viewpoint planning on a calibrated color checkerboard, a Triangular sandstone slab, and a Gypsum boulder with Iron Oxide deposits as seen in Fig.  5. The color checker is a highly diverse surface of tiled color squares with varying albedo. The rocks used in this experiment are notional examples for complex geometries derived from inspection and identification of target minerals. For each viewpoint, if planning is not feasible for either the arm or Stewart platform due potential collisions or unreachable poses beyond the operable workspace, we discard that viewpoint and continue.
@@ VII-B Point Alignment
From the results in Section IV, we know PROSPECT has minimal pose error. Building on these results, we design two experiments to explore the precision of the device. Using the color checker from the previous experiment, we plan a trajectory to directly over a red-colored square at the nominal working distance. After a sample is acquired, the arm is linearly translated with PROSPECT in the same orientation along the upwards $z$ axis by 12 mm. The arm is then perturbed by adding a small angular offset to the rotation in two increments of 11 and 17 degrees. This set of experiments is designed to highlight small errors induced by sample-based motion planning methods.
@@ VIII Results & Discussion
The results of the probe alignment underscore the need for precision alignment with fiber optics. Fig. 6 shows the reflectance profiles collected under each of the scenarios. The nominal reflectance profile shows a markedly different shape from each of the other test scenarios. In previous research [30] we held a uniformity assumption regarding the composition of the scanned surface and expected the spectral signal magnitude to vary linearly as a function of distance to the surface. As modeled through Algorithm 2, as distance increases the number of points contributing to the spectral signature likewise increases, but at a quadratic rate. The increased scan height results in adjacent color cells contributing to observed reflectance profile. Similarly when the probe is rotated out of alignment by small angles, the conic intersection occurs at an oblique angle, with a larger spread of points beyond the intended intersection area.
Figure 6: Reflectance P=profiles for different probe alignments with fixed $T_{stew,xy}$
Figure 7: Point cloud of gypsum boulder with sampled viewpoints and corresponding normals, colorized by their SAM score compared to PROSPECT using the method proposed in Algorithm  1. (a) Compared to viewing points directly from above with no rotation. (b) Compared to where Stewart platform always points straight down, which may not be parallel to the surface normal.
@@
While Fig. 6 demonstrates the errors induced by small misalignments in motion planning, we also quantitatively justify the use of PROSPECT to match surface normals. We select two alternative approaches to demonstrate the effectiveness of our end-effector: one where all normals in Algorithm  1 are overwritten to be parallel to Unit Vector, and one where in Algorithm  1 all $\mathbf{R_{stew}}$ are replaced with unit rotation. The first alternative approach is designed to emulate viewpoint planning for PROSPECT on a 3-axis gantry, and the second alternative approach is designed to emulate poor alignment of the spectrometer relative to an object’s surface normal. As demonstrated in Fig.  5, orienting the spectrometer to be parallel to the normal of the surface yields the strongest signal amplitude, and as such is assumed to be optimal for downstream processing. The alternative approaches’ spectral signature differ from that of the normal matching baseline, suggesting that spectral profiling quality would suffer if not for the use of PROSPECT. We calculate dissimilarity between any two collected point clouds through the Spectral Angle Mapper (SAM), with results shown in Fig.  6.
@@
SAM is well supported by spectroscopy literature as a means to calculate the divergence between two spectral signals as a similarity metric [40] while showing robustness to illumination intensity. In this formulation $s_{p}$ is the spectral vector observed using PROSPECT, and $\hat{s}$ represents the correlated spectra vector collected through one of the ablated motion planning routines. SAM is defined over $[0,\pi]$, where 00 indicates a high degree of similarity and $\pi$ high dissimilarity.
Fig. 6 shows the dissimilarity in the acquired measurements. Fig. 6a presents the results of alignment along the unit vector compared to PROSPECT. Table I highlights the mean, standard deviation and range of the observed values. Both strategies indicate a wide range of SAM scores, indicating spectral measurements do not uniformly diverge across the surface. In both subfigures of Fig. 7, there is a corresponding decrease in similarity as the normal diverges from alignment with the $z$ axis. Points on a relatively flat surface exhibit close agreement with measurements scanned without the spherical alignment, as they are not necessary in these scenarios.
@@ IX Conclusions
In this work, we presented a first of its kind integrated end-effector – PROSPECT – for the precise control of a fiber optic cable for automated spectral measurements of complex geometries. Our proposed motion planning and physics-based point attribution allows for the progressive construction of spectral-spatial point clouds.. We demonstrated the validity of our approach in increasingly complex scenarios with variable spectral profiles. For generality to the robotics community, the system is designed to be agnostic to the model of robot and spectrometer.
In future works, we plan to expand our approach of scanning objects to identify and segment non-visible material imperfections through the use of automated anomaly detection, as in our previous work. Additionally, we will increase the speed of the motion planning through coupled motion planning of the end-effector and PROSPECT for smooth continuous trajectories over object contours. We believe this work represents a practical step towards the fusion of spectral sciences and robotics, and opens new opportunities to non-destructive characterization in manufacturing and automation.

### v2

@@ V-C Surface Normal Matching
To orient the fiber optic normal to the surface at a fixed depth, we implement a surface normalization algorithm. Given the initial pose of the Stewart platform, the positioning and readings of three ToF sensors onboard the end effector, and a target offset distance, a new pose for the platform is computed. The platform is then commanded to move to this position using the joint angles computed by the inverse kinematics algorithm.
@@ VI Spectral-Spatial Modeling
Despite their slender form, fiber optic cables acquire light from the environment in a more complex manner than tracing a ray from the fiber terminus to the surface. We begin by modeling the interaction of light reflected from a target object and into the ferrule of the fiber optic cable. Numerical aperture (NA) defines the maximum angle from which light will be accepted into the fiber and is given by: $NA=n_{0}\sin{\theta_{max}}$ where $n_{0}$ defines the refractive index of the medium through which light transits before entering the fiber, which we assume to be air with a refractive index of 1.0. $\theta_{max}=\arcsin{NA}$ yields the half-angle forming the acceptance cone.
Fig. 4 shows the fiber optic cable above a nominal surface $S$ that the PROSPECT platform will scan. The acceptance cone is centered on the ferrule. The distance from the tip of the ferrule to the surface is given as $d$, which intersects with the surface in a circle with a radius $r$. $d$ is approximated from a static 3D scan of the object surface and confirmed using the 3 ToF distance sensors. The size of the cone base varies as a function of $d$, but $\theta_{max}$ remains constant. The area of the circular intersection region is given as $A=\pi(d\tan{(\arcsin{(NA)}}))^{2}$.
Figure 4: Model of spectroscopic measurement and modeling of object reflectance. Light is emitted from the PROSPECT platform (yellow) and reflected (green) by the surface $S$ (blue). The conical acceptance profile is defined by offset distance from the surface $d$, subsurface penetration $\epsilon$, and acceptance angle $\theta_{max}$. The intersection of the acceptance cone with the point cloud yields a subset of points $S_{c}$ (pink) with which the observed spectral signature is associated.
@@
To mark points as having been explored, we find the intersection between the surface $S$ and the circular profile of the acceptance cone projected on the surface, $S_{c}$. The intersection is found by creating a 3D mesh of the acceptance cone projected slightly below the surface of $S$, at a distance $\epsilon$. The mesh is then transformed into a Delaunay triangulation, which is queried to find points from $S$ that lie inside it, $S_{c}$. Since scanning occurs over the surface at a close distance, the number of points observed follows the following relationship: $1\leq||S_{c}||<<||S||$. As the end-effector is translated according to Algorithm 1, the point cloud is annotated so that each point $x_{i},y_{i},z_{i}$ in $S_{c}$ is associated with a spectral vector $[\lambda_{0},\lambda_{1},\dots,\lambda_{n}]$.
As the size of $A$ increases, the intersecting area and points in $S_{c}$ also increase. However, as the distance $d$ increases, the number of points contributing to the spectral signal similarly increases, while the overall intensity of the signal decreases because fewer photons enter the fiber probe. This mixing of spectral signatures from multiple points complicates the interpretation and classification of the signal. Knowing the approximate point density of the point cloud $S$, we chose a constant value for $d$ to avoid variable signal quality. Algorithm 2 details the steps to extract the points $S_{c}$ and form the correlated spectral-spatial measurements.
@@ VII-A Sparse Point Acquisition
The first set of experiments demonstrates the creation of a sparse spectral-spatial point cloud. These scans are designed to be completed in less than 5 minutes to broadly understand an object’s surface composition. Fig. 5 shows the variation of spectral signatures across the surface of an object and emphasizes the importance of collecting precise spectral measurements. In each of these scenarios, viewpoints were planned with a voxel size of 5 cm.
We test our viewpoint planning on a calibrated color checkerboard, a triangular sandstone slab, and a Gypsum boulder with Iron Oxide deposits as seen in Fig. 5. The color checker is a highly diverse surface of tiled color squares with varying albedo. The rocks used in this experiment are notional examples of complex geometries derived from inspection and identification of target minerals. For each viewpoint, if planning is not feasible for either the arm or Stewart platform due to potential collisions or unreachable poses beyond the operable workspace, we discard that viewpoint and continue.
@@ VII-B Point Alignment
From the results in Section IV, we know PROSPECT has minimal pose error. Building on these results, we design two experiments to explore the precision of the device. Using the color checker from the previous experiment, we plan a trajectory to directly over a red-colored square at the nominal working distance. After acquiring a sample, the arm is linearly translated with PROSPECT in the same orientation along the upward $z$ axis by 12 mm. The arm is then perturbed by adding a small angular offset to the rotation in two increments of 11 and 17 degrees. This set of experiments is designed to highlight small errors induced by sample-based motion planning methods.
@@ VIII Results & Discussion
The results of the probe alignment underscore the need for precision alignment with fiber optics. Fig. 6 shows the reflectance profiles collected in each of the scenarios. The nominal reflectance profile shows a markedly different shape from each of the other test scenarios. In previous research [29], we assumed a uniformity in the composition of the scanned surface and expected the magnitude of the spectral signal to vary linearly as a function of distance from the surface. As modeled through Algorithm 2, as the distance increases, the number of points contributing to the spectral signature also increases, but at a quadratic rate. The increased scan height results in adjacent color cells contributing to the observed reflectance profile. Similarly, when the probe is rotated out of alignment by small angles, the conic intersection occurs at an oblique angle, with a larger spread of points beyond the intended intersection area.
Figure 6: Reflectance profiles for different probe alignments with fixed $T_{stew,xy}$.
Figure 7: Point cloud of gypsum boulder with sampled viewpoints and corresponding normals, colorized by their SAM score compared to PROSPECT using the method proposed in Algorithm 1. (a) Compared to viewing points directly from above with no rotation. (b) Compared to where Stewart platform always points straight down, which may not be parallel to the surface normal.
@@
While Fig. 6 demonstrates the errors induced by small misalignments in motion planning, we also quantitatively justify the use of PROSPECT to match surface normals over a complex geometry. We select two alternative approaches to demonstrate the effectiveness of our end-effector: one where all normal vectors in Algorithm 1 are overwritten to be parallel to the unit vector, and one where in Algorithm 1 all $\mathbf{R_{stew}}$ are replaced with unit rotation. The first alternative approach is designed to emulate viewpoint planning for PROSPECT on a 3-axis gantry, and the second alternative approach is designed to emulate poor alignment of the spectrometer relative to an object’s surface normal. As demonstrated in Fig. 6, orienting the spectrometer to be parallel to the normal of the surface yields the strongest signal amplitude and as such is assumed to be optimal for downstream processing. The alternative approaches’ spectral signatures differ from that of the normal matching baseline, suggesting that spectral profiling quality suffers without PROSPECT. We calculate the dissimilarity between any two collected point clouds using the Spectral Angle Mapper (SAM), with results shown in Fig. 7.
@@
SAM is well supported in spectroscopy literature as a means to calculate the divergence between two spectral signals as a similarity metric [40] while showing robustness to illumination intensity. In this formulation $s_{p}$ is the spectral vector observed using PROSPECT, and $\hat{s}$ represents the correlated spectra vector collected through one of the ablated motion planning routines. SAM is defined over $[0,\pi]$, where 00 indicates a high degree of similarity and $\pi$ a high dissimilarity.
Fig. 7 shows the dissimilarity in the acquired measurements. Fig. 7a presents the results of alignment along the unit vector compared to PROSPECT. Table I highlights the mean, standard deviation, and range of the observed values. Both strategies exhibit a wide range of SAM scores, indicating that spectral measurements do not diverge uniformly across the surface. In both subfigures of Fig. 7, there is a corresponding decrease in similarity as the normal diverges from alignment with the $z$ axis. Points on a relatively flat surface exhibit close agreement with measurements scanned without spherical alignment since the platform was already neutrally aligned in those scenarios.
@@ IX Conclusions
In this work, we presented PROSPECT, an integrated end-effector for the precise control of a fiber optic cable for automated spectral measurements of complex geometries. Our proposed motion planning and physics-based point attribution allows for the progressive construction of spectral-spatial point clouds. We demonstrated the validity of our approach in increasingly complex scenarios with variable spectral profiles. For generality to the robotics community, the system is designed to be agnostic to the robot and spectrometer model.
In future works, we plan to expand our approach of scanning objects to identify and segment non-visible material imperfections through the use of automated anomaly detection, as in our previous work [41]. New ToF sensors offer multizone ranging; allowing for the three ToF sensors to be condensed to a single unit and streamline the electronics in the top platform. Additionally, we will increase the speed of motion planning through coupled motion planning of the end-effector and PROSPECT for smooth continuous trajectories over object contours. We believe this work represents a practical step towards the fusion of spectral sciences and robotics and opens new opportunities to non-destructive characterization in manufacturing and automation.